# Binary Stars Approaching Supermassive Black Holes: Tidal Break-up, Double Stellar Disruptions and Stellar Collision

### v1

@@ 6.2.2 Merger Remnants from Stellar Collsions
(Note that $\epsilon_{\rm rem}$ can be much larger than unity; see Figure 8.) Comparing to Equation (28), we see that typically $P_{\rm rem}\gg P_{\rm bs}$. Thus these merger remnants could also provide an explanation for the repeated partial TDEs, especially for smaller $M/m$ or/and low-mass stars. There are similar difficulties in this scenario compared to the bound star scenario (see Section 6.2.1). One advantage of the merger remnant scenario is that because of dissipation in the merger process, the merger remnant stars likely have a very extended envelope. This makes their partial disruptions more likely even without significant orbital decay due to gravitational radiation.

### v2

@@ 6.2.2 Merger Remnants from Stellar Collsions
(Note that $\epsilon_{\rm rem}$ can be much larger than unity; see Figure 11.) Comparing to Equation (29), we see that typically $P_{\rm rem}\gg P_{\rm bs}$. Thus these merger remnants could also provide an explanation for the repeated partial TDEs, especially for smaller $M/m$ or/and low-mass stars. One advantage of the merger remnant scenario is that because of dissipation in the merger process, the merger remnant stars likely have a very extended envelope. This makes their partial disruptions more likely even without significant orbital decay due to gravitational radiation. Bradnick et al. (2017) have suggested that stellar mergers can produce strong magnetic fields, and the disruption of such stars may lead to jetted TDEs.
@@
We are grateful to Re’em Sari for useful discussions, and the anonymous referee for constructive comments and suggestions that greatly improved the manuscript.